a pendulum bob of mass 80mg and carring a charge of 0.00000002C is at rest in auniform horizontalk charge electric field of 20kVm^-1 find tension in thread

Grade:12

1 Answers

Arun
4 years ago
As the electric field is in horizontal direction and the bob is positively charged it will move in the direction of the field. Lets define the angle made by the bob after displacement from y-axis as alpha from the y axis. Resolving the components we have.Tsin alpha = qE 1Tcos alpha = mg 2Dividing 1by 2 , we have Tanalpha = qE/mg Mass= 80 * 10^-6 kg. g = 9.8m/sq= 2*10^-8. E= 2*10000 Solving the above equation we have Tan alpha = .5012 (approx)Alpha = 27degrees Now from 1T= mg/sinalpha putting the values and solvingT = 8.8*10^-4 newton(approximately) Hope it helps..cheers